mathslover (mathslover):

Convert into polar form : (-1/2) + i ( \(\frac{-\sqrt{3}}{2}\) )

mathslover (mathslover):

(-1/2) + i ( \(-\frac{\sqrt{3}}{2}\) )

mathslover (mathslover):

I got : cos ( 4 pi/3 ) + sin ( 4 pi/3)

mathslover (mathslover):

But the book says : cos ( - 2pi/3 ) + i sin ( - 2pi/3). Which is also right...

OpenStudy (anonymous):

isin

OpenStudy (anonymous):

You're wrong.

mathslover (mathslover):

Sorry typing mistake I meant i sin 4pi/3

mathslover (mathslover):

^ cos ( 4 pi/3 ) + i sin(4pi/3)

OpenStudy (anonymous):

Find the absolute value of this: \[\left| -\frac{ 1 }{ 2 } -\frac{ i \sqrt{3} }{ 2 }\right|\]

mathslover (mathslover):

So which should I prefer more ? -2pi/3 or 4 pi/3

OpenStudy (anonymous):

So \[=\sqrt{(-\frac{ 1 }{ 2 })^2+(-\frac{ \sqrt{3} }{ 2 })^2}\]

OpenStudy (anonymous):

So r=1

OpenStudy (anonymous):

Now draw it.

OpenStudy (anonymous):

Drawing it gives you a way better picture of what the angle is.

mathslover (mathslover):

\[\large{\sqrt{ (\frac{-1}{2} )^2 + ( \frac {- \sqrt{3}}{2})^2}}\] right so it is : \(\1\)

OpenStudy (anonymous):